3r-2s=5 and 2r+3s=12
Answer by mananth(16949)

3 r -2 s = 5 .............1
2 r + 3 s = 12 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
9 r -6 s = 15
4 r 6 s = 24
Add the two equations
13 r = 39
/ 13
r = 3
plug value of r in (1)
3 r + -2 s = 5
9 + -2 s = 5
-2 s = 5 -9
-2 s = -4
s = 2